OpenStudy (anonymous):

What percentage interest rate is a loan that charges one hundred dollars in simple annual interest over 5 years on a one thousand dollar balance? zero point two four percent two percent twenty percent twenty-four percent

OpenStudy (anonymous):

@Nurali posted :)

OpenStudy (nurali):

I = PRT 100 = 1,000 * R * 5 100 = 5,000R 100/5,000 = R 0.02 = R Rate = 2%

OpenStudy (anonymous):

Nadia took two practice exams to prepare for her college entrance exam. On the first practice exam, she scored one thousand two hundred. After studying, her score increased by twenty percent. What was her score on the second practice exam? one thousand two hundred twenty one thousand four hundred one thousand four hundred forty one thousand five hundred

OpenStudy (anonymous):

8 more questions :D

OpenStudy (nurali):

one thousand two hundred twenty

OpenStudy (anonymous):

Fayed tipped the pizza delivery driver three dollars for pizzas that cost twenty-four dollars. What percent did he tip the driver? three percent eight percent twelve point five percent twenty-one percent

OpenStudy (nurali):

twenty-one percent

OpenStudy (anonymous):

Which of the following forms a proportion with nine over fifteen? twelve over twenty ten over six fifteen over twenty-one three over forty-five

OpenStudy (anonymous):

7 more questions.

OpenStudy (nurali):

ten over six don't know

OpenStudy (anonymous):

thats fine

OpenStudy (anonymous):

Danika has saved enough money to pay cash for a used car. She wants to buy a car that costs four thousand nine hundred dollars before taxes. If she will be charged seven percent sales tax, what is the total amount she will pay for the car? four thousand five hundred fifty-seven dollars four thousand nine hundred seventy dollars five thousand two hundred forty-three dollars five thousand six hundred dollars

OpenStudy (nurali):

four thousand five hundred fifty-seven dollars

OpenStudy (anonymous):

Which of the following would you use as the percent equation to determine what number is thirty percent of ninety? a equals zero point three times ninety ninety equals zero point three times a a equals zero point nine times thirty thirty equals zero point nine times a

OpenStudy (anonymous):

@Nurali

OpenStudy (nurali):

thirty equals zero point nine times a don't know.

OpenStudy (anonymous):

Lydia received a parking ticket because she parked ten feet from a fire hydrant. She had estimated her distance from the fire hydrant as being the minimum allowable distance, which is fifteen feet. What was her percent error? fifteen percent thirty-three percent fifty percent sixty-seven percent

OpenStudy (anonymous):

3 more!!!!!!!

OpenStudy (nurali):

fifty percent

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What percentage interest rate is a loan that charges in simple annual interest over 5 years on a balance?

OpenStudy (nurali):

@shyshy_falcon3113 Principal = P = $5000 Simple interest = I = $5000 Rate percent p.a. = r Period = t = 5 yr I = (P r t)/100 => 5000 = (5000 * r * 5)/100 => 100 = 5r => r = 100/5 = 20 Answer